Voltage of 2 volt exist across a potentiometer wire of length 4m when the P.O. across 2 resistance of second circuit is measured by this potentiometer , the balancing length is found to be 4cm . The current in the second circuit is ?

Grade:11

1 Answers

Khimraj
7 years ago
In potentiometer the votage drop across wire in primary circuit is balanced by the potential drop of secondary circuit. So current in secodandary circuit is given by (balancing voltage)/resistanceBalancing voltage = x*l, x = voltage per unit length = 2/4 = 0.5Balancing voltage = 0.5*0.04 = 0.02 Current = 0.02/2 = 0.01